Your piggy bank contains $19.75 in dimes and quarters. There are 100 coins in all. How many dimes are there?
A. 25
B. 30
C. 35
D. 40

Answer by jim_thompson5910(35256)

d+q = 100 ---> q = 100-d

0.1d+0.25q=19.75

10d+25q=1975

10d+25(100-d)=1975

10d+2500-25d=1975

-15d+2500=1975

-15d=1975-2500

-15d=-525

d=(-525)/(-15)

d=35


So there are 35 dimes.
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------